OpenStudy (anonymous):

BEEN STUCK ON THIS ONE FOREVER HELP PLEASE?? the mean value theorem applies to f(x)=1/x on the interval [1,3],find all values of c in the interval such that f'(c)=f(3)-f(1)/3-1

OpenStudy (anonymous):

i finally got it !!

OpenStudy (anonymous):

take the derivative of f(x), which should be f'(x) = -1/x^2. Set that expression equal to whatever f(3)-f(1)/3-1 is. Then solve.

OpenStudy (anonymous):

f'(x) = -1/x^2 By MVT, f'(c) = -1/c^2 = (1/3 - 1/1)/(3-1) = (1-3)/6 = -1/3 So, c = +/-sqrt(3)

OpenStudy (anonymous):

congratulation

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

yes !! lol been on it for 2hrs